Given two commodities A and B, where price of A is ₦50 and the price of B is ₦60. If the consumer’s income is ₦1200 and the utility function U = AB, Find: i. the equilibrium values of the two commodities. ii. Mention four properties of the demand function derived in (i) above


1
Expert's answer
2021-03-01T11:31:54-0500

Solution:

i). Equilibrium values of the two commodities:

PA = 50

PB = 60

I = 1200

U (A, B) = AB

Derive Budget line:

I = PAA + PBB

1200 = 50A + 60B


"\frac{MU_{A} }{MU_{B}}" = "\frac{P_{A} }{P_{B}}"


U (A, B) = AB

MUA = "\frac{\partial _{u} }{\partial _{A}}" = B


MUB = "\frac{\partial _{u} }{\partial _{B}}" = A


"\frac{MU_{A} }{MU_{B}}" = "\frac{B }{A}"


Therefore: "\frac{MU_{A} }{MU_{B}}" = "\frac{P_{A} }{P_{B}}"


"\frac{B }{A}" = "\frac{50}{60}"


B = "\frac{5 }{6} A"

Solve for A:

I = PAA + PBB

1200 = 50A + 60B

1200 = 50A + 60 "(\frac{5 }{6} A)"

1200 = 50A + 50A

1200 = 100A

A = 12

Derive B:

B = "\frac{5 }{6} A"


B ="(\frac{5 }{6}) 12"


B = 10

U (A, B) = AB

U (12, 10) = "12\times 10 = 1,200"


The equilibrium value of commodity A = 12

The equilibrium value of commodity B = 10


ii). The four properties of the demand function derived above include the following:

·        The demand function is single-valued.

·        The demand function is insensitive to proportional increases in price and income.

·        The demand function exists.

·        The demand function exhausts the consumer’s budget.